a-10 over 3 = -4. solve for a.
Answer by checkley77(12569)

a-10 over 3 = -4
(a-10)/3=-4
a-10=3*-4
a-10=-12
a=-12+10
a=-2 ans.
OR:
a-10/3=-4
a=-4+10/3
a=(-12+10)/3
a=-2/3